# Split Two-Periodic Aztec Diamond

### v4

@@ 3. Derivation of the Correlation Kernel
Our main goal in this section is to prove the results stated in Theorem 2.1. In order to prove Theorem 2.1, we start in Section 3.1 by stating an intermediate version of the kernel which can be derived by using techniques from [3]. The form of the kernel differs slightly depending on whether the coordinate $(4m^{\prime},2\xi^{\prime}+j)$ is on the $\alpha$-side or $\beta$-side of the interface. In Section 3.2, we prove the statement of the kernel for $m^{\prime}\leq\nicefrac{{N}}{{2}}$ and in Section 3.3, we prove the statement of the kernel for $m^{\prime}>\nicefrac{{N}}{{2}}$.
@@ 3.1. An Intermediate Correlation Kernel
where the function $g_{\alpha,\beta}(z)$ is defined in equation (14). Since $\det\phi_{N}(z)=1$, the eigenvalues of $\phi_{N}(z)$ can be expressed as,
@@
Let $0<m<N$, and $-N\leq\xi,\xi^{\prime}\leq-1$. The split two-periodic Aztec diamond of size $2N$ has a correlation kernel given by,
@@
$\phi_{N}(w)=\Big{(}r_{\alpha,1}(w)^{\frac{N}{2}}F_{\alpha,1}(w)+r_{\alpha,2}(w)^{\frac{N}{2}}F_{\alpha,2}(w)\Big{)}\Big{(}r_{\beta,1}(w)^{\frac{N}{2}}F_{\beta,1}(w)+r_{\beta,2}(w)^{\frac{N}{2}}F_{\beta,2}(w)\Big{)}$
@@
(39) $F_{\alpha,k_{1}}(w)\phi_{N}(w)=r_{\alpha,k_{1}}(w)^{\frac{N}{2}}F_{\alpha,k_{1}}(w)\Big{(}r_{\beta,1}(w)^{\frac{N}{2}}F_{\beta,1}(w)+r_{\beta,2}(w)^{\frac{N}{2}}F_{\beta,2}(w)\Big{)}$
Combining the above with equation (38) gives,
@@
$\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,k_{1}}(w)^{-m^{\prime}}r_{\alpha,k_{2}}(w)^{\frac{N}{2}}F_{\alpha,k_{1}}(w)F_{1,N}(w)F_{\alpha,k_{2}}(w)F_{\varepsilon,k}(z)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
@@
For any $\varepsilon\in(0,1)$, the matrix valued functions $F_{\varepsilon,1}(z)$ and $F_{\varepsilon,2}(z)$ are analytic for all $z\in\mathbb{C}\backslash\{(-\infty,-\varepsilon^{-2}]\cup[-\varepsilon^{2},0]\}$.
@@
(44) $\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,1}(w)^{\frac{N}{2}-m^{\prime}}F_{\alpha,1}(w)F_{1,N}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ =\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{2,N}(w)r_{\alpha,1}(w)^{N-m^{\prime}}r_{\beta,1}(w)^{\frac{N}{2}}\\ \times F_{\alpha,1}(w)F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
@@
(51) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\alpha}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,1}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\$
@@
(52) $\frac{2F_{\alpha,1}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g(w)}=F_{\alpha,1}(w)$
@@
(53) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\alpha}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k_{1}=1,2}\sum_{k_{2}=1,2}\sum_{k_{3}=1,2}\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,k_{1}}(w)^{\frac{N}{2}-m^{\prime}}r_{\alpha,k_{2}}(w)^{-\frac{N}{2}}r_{\alpha,k_{3}}(w)^{\frac{N}{2}}\\ \times F_{\beta,k_{1}}(w)F_{\alpha,k_{2}}(w)F_{1,N}(w)F_{\alpha,k_{3}}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
@@
$k_{3}=1$ pole of order $<N$ pole of order $<2N$
$k_{3}=2$ zero of order $>N$ zero of order $>0$
@@
$k_{3}=1$ pole of order $<3N$ pole of order $<4N$
$k_{3}=2$ pole of order $<N$ pole of order $<2N$
@@
(54) $\sum_{k_{1}=1,2}\sum_{k_{2}=1,2}\sum_{k_{3}=1,2}\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,k_{1}}(w)^{\frac{N}{2}-m^{\prime}}r_{\alpha,k_{2}}(w)^{-\frac{N}{2}}r_{\alpha,k_{3}}(w)^{\frac{N}{2}}\\ \times F_{\beta,k_{1}}(w)F_{\alpha,k_{2}}(w)F_{1,N}(w)F_{\alpha,k_{3}}(z)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ =\sum_{k=1,2}\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,k}(w)^{\frac{N}{2}-m^{\prime}}r_{2,N}(w)r_{\alpha,1}(w)^{\frac{N}{2}}r_{\beta,1}(w)^{\frac{N}{2}}\\ \times F_{\beta,k}(w)F_{\alpha,1}(w)F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\sum_{k=1,2}\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,k}(w)^{\frac{N}{2}-m^{\prime}}r_{\alpha,1}(w)^{\frac{N}{2}}r_{2,N}(w)r_{\beta,1}(w)^{\frac{N}{2}}\\ \times F_{\beta,k}(w)F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\$
@@
(55) $\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{N-m^{\prime}}r_{2,N}(w)r_{\alpha,1}(w)^{\frac{N}{2}}F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\$
@@
(56) $\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
@@
$\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,1}(w)F_{\alpha,1}(w)=F_{\alpha,1}(w)-\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\alpha,1}(w)F_{\beta,2}(w)F_{\alpha,2}(w)$
@@ 4.1. Preliminaries of Saddle Functions
We know that $\psi_{k}(z)$ always has at least one saddle point between the branch cuts on the negative real axis. We denote this saddle point $z_{k}^{*}$. [^8] This is the same notation used in [11]. We start by stating,
@@
This is largely a subset of Lemma 6.8 in [11]. There are a few key differences we take the time to acknowledge and justify. Firstly, since there is an overall sign change between their saddle function and the saddle function presented here, there is a reversal of maxima and minima. Additionally, while the asymptotic coordinate is restricted to the smooth region in the proof presented by Duits and Kuijlaars, we address cases where this is not necessarily true. We present the proof of these additional part below.
@@
Since we are restricting our analysis to $x<\nicefrac{{1}}{{2}}$ we can make some additional remarks regarding the general location of the saddle points. Under this restriction, $\psi_{1}(z)$ always has a single saddle point, while $\psi_{2}(z)$ has three saddle points (unless the point $(x,y)$ is on a macroscopic boundary). We will always refer to the other two saddle points of $\psi_{2}(z)$ as $z_{3}^{*}$ and $z_{4}^{*}$. When $(x,y)\in\mathfrak{S}$, these additional saddle points also lie on the interval $(-\alpha^{-2},-\alpha^{2})$. We adopt the convention, $z_{3}^{*}<z_{2}^{*}<z_{4}^{*}$. The points $z_{3}^{*}$ and $z_{4}^{*}$ are both local minimums of $\operatorname{Re}\psi_{2}(z)$.
The contour of steepest ascent from $z_{k}^{*}$ is the contour through $z_{k}^{*}$ where $\operatorname{Im}\psi_{k}(z)$ is held constant and $\operatorname{Re}\psi_{k}(z)$ increases the fastest as we move away from the saddle point. Since $z_{1}^{*}$ is a maximum of $\operatorname{Re}\psi_{1}(z)$, the path of steepest ascent is perpendicular to the real axis at $z_{1}^{*}$ (the same is true for $z_{2}^{*}$ when the asymptotic coordinate is in the smooth region). We define the following regions of the complex plane, [^9] While this notation is the same as in [11] the regions will be reversed in our notation due to the change in sign of the saddle functions presented here.
@@
    The steepest ascent contour of $\psi_{1}(z;x,y)$ through $z_{1}^{*}$, denoted $\gamma_{s,1}$, is a simple closed curve that contains the interval $[-\alpha^{2},0]$ and intersects the positive real line at $z=1$.
@@
    $\Omega^{+}_{1}$ is a bounded set containing a single component which contains $\gamma_{s,1}\backslash\{z_{1}^{*}\}$.
@@
    $\Omega^{-}_{1}$ is an open set with an unbounded component that contains a contour going around $(-\infty,-\alpha^{-2}]$ and with a bounded component that contains a contour around $[-\alpha^{2},0]$.
@@
    The steepest ascent contour of $\psi_{2}(z;x,y)$ through $z_{2}^{*}$, denoted $\gamma_{s,2}$, is a simple closed curve that contains the interval $[-\alpha^{2},0]$ and intersects the positive real line at $z=1$.
@@
    $\Omega^{+}_{2}$ is a bounded set containing at most three components that are a positive distance apart. One of which contains $\gamma_{s,2}\backslash\{z_{2}^{*}\}$.
@@
    $\Omega^{-}_{2}$ is an open set with an unbounded component that contains a contour going around $(-\infty,-\alpha^{-2}]$ and with a bounded component that contains a contour around $[-\alpha^{2},0]$.
Figure 9 shows an example of the regions $\Omega_{1}^{+}$ and $\Omega_{1}^{-}$ along with the steepest ascent contour. This lemma is a slightly modified version of Lemma 6.9 of [11]. The reader should refer to their text for a the proof. One can check that the proof is valid for $\psi_{1}(z)$ for any $x<\nicefrac{{1}}{{2}}$. This is because $z_{1}^{*}$ is always a local maximum of $\operatorname{Re}\psi_{1}(z)$. In the first part, we conclude that $\Omega^{+}_{1}$ is a single component, because there is exactly one saddle point on the interval $(-\alpha^{-2},-\alpha^{2})$. The case of three components is only possible if there are three saddle points on the interval $(-\alpha^{-2},-\alpha^{2})$.
@@
[…]
@@
The above information is enough for the analysis of $I_{2,1}^{\alpha}$ and for the analysis of $I^{\alpha}_{2,2}$ in the smooth region. In order to analyze $I^{\alpha}_{2,2}$ in the rough region, we will need some information on the contours of steepest descent and ascent in the rough region. For $x\in(0,\nicefrac{{1}}{{2}})$ and $(x,y)\in\mathfrak{R}$, the additional saddle points of $\psi_{2}9z)$, $z_{3}^{*}$ and $z_{4}^{*}$, are complex. In particular, $z_{3}^{*}=\overline{z_{4}^{*}}$. We will let $z_{3}^{*}$ denote the complex saddle point with positive imaginary part. We define the following regions of the complex plane,
@@
    $\Pi^{+}$ contains at least two and at most three bounded components. One component will always contain $z=1$, while the other guaranteed component contains at least one of the branch points, $-\alpha^{2}$ or $-\alpha^{-2}$. The third component, if it exists, must contain the other branch point, intersect the branch cut, and be positive distance away from the other components.

### v5

@@ 3. Derivation of the Correlation Kernel
Our main goal in this section is to prove the results stated in Theorem 2.1. In order to prove Theorem 2.1, we start in Section 3.1 by stating an intermediate version of the kernel which can be derived by using techniques from [BD19]. The form of the kernel differs slightly depending on whether the coordinate $(4m^{\prime},2\xi^{\prime}+j)$ is on the $\alpha$-side or $\beta$-side of the interface. In Section 3.2, we prove the statement of the kernel for $m^{\prime}\leq\nicefrac{{N}}{{2}}$ and in Section 3.3, we prove the statement of the kernel for $m^{\prime}>\nicefrac{{N}}{{2}}$.
@@ 3.1. An Intermediate Correlation Kernel
which follows from the formulas in (9), (10), (11). Since $\det\phi_{N}(z)=1$, the eigenvalues of $\phi_{N}(z)$ can be expressed as,
@@
Let $0<m<N$, and $N\leq\xi,\xi^{\prime}\leq-1$. The split two-periodic Aztec diamond of size $2N$ has a correlation kernel given by,
@@
$\phi_{N}(w)=\Big(r_{\alpha,1}(w)^{\frac{N}{2}}F_{\alpha,1}(w)+r_{\alpha,2}(w)^{\frac{N}{2}}F_{\alpha,2}(w)\Big)\Big(r_{\beta,1}(w)^{\frac{N}{2}}F_{\beta,1}(w)+r_{\beta,2}(w)^{\frac{N}{2}}F_{\beta,2}(w)\Big).$
@@
(37) $F_{\alpha,k_{1}}(w)\phi_{N}(w)=r_{\alpha,k_{1}}(w)^{\frac{N}{2}}F_{\alpha,k_{1}}(w)\Big(r_{\beta,1}(w)^{\frac{N}{2}}F_{\beta,1}(w)+r_{\beta,2}(w)^{\frac{N}{2}}F_{\beta,2}(w)\Big).$
Above we used the fact that $F_{\alpha,1}(z)^{2}=F_{\alpha,1}(z)$ and $F_{\alpha,1}(z)F_{\alpha,2}(z)=0$. Combining the above with equation (36) gives,
@@
$\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}z\text{d}w}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,k_{1}}(w)^{-m^{\prime}}r_{\alpha,k_{2}}(w)^{\frac{N}{2}}F_{\alpha,k_{1}}(w)F_{1,N}(w)F_{\alpha,k_{2}}(w)F_{\varepsilon,k}(z)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
@@
[…]
@@
(42) $\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}z\text{d}w}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,1}(w)^{\frac{N}{2}-m^{\prime}}F_{\alpha,1}(w)F_{1,N}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ =\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}z\text{d}w}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{2,N}(w)r_{\alpha,1}(w)^{N-m^{\prime}}r_{\beta,1}(w)^{\frac{N}{2}}\\ \times F_{\alpha,1}(w)F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
@@
(49) $\Big[\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big]_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\alpha}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,1}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}.$
@@
(50) $\frac{2F_{\alpha,1}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g(w)}=F_{\alpha,1}(w).$
One can check the above equality using the explicit formulas given in equations (9), (10), (11).
Remark 3.1.
We are not surprised that an equality like this. Since the $F$-matrices are rank one, we expect the product $F_{\alpha,1}(w)F_{\beta,1}(w)F_{\alpha,1}(w)$, to be a scalar multiple of $F_{\alpha,1}(w)$. Why does our integral produce the “right” constant for simplification? I do not have a satisfactory answer to this question.
@@
(51) $\Big[\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big]_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\alpha}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k_{1}=1,2}\sum_{k_{2}=1,2}\sum_{k_{3}=1,2}\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}z\text{d}w}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,k_{1}}(w)^{\frac{N}{2}-m^{\prime}}r_{\alpha,k_{2}}(w)^{-\frac{N}{2}}r_{\alpha,k_{3}}(w)^{\frac{N}{2}}\\ \times F_{\beta,k_{1}}(w)F_{\alpha,k_{2}}(w)F_{1,N}(w)F_{\alpha,k_{3}}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
@@
$k_{3}=1$ pole of order <N<N pole of order <2N<2N
$k_{3}=2$ zero of order >N>N zero of order >0>0
@@
$k_{3}=1$ pole of order <3N<3N pole of order <4N<4N
$k_{3}=2$ pole of order <N<N pole of order <2N<2N
@@
(52) $\sum_{k_{1}=1,2}\sum_{k_{2}=1,2}\sum_{k_{3}=1,2}\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}z\text{d}w}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,k_{1}}(w)^{\frac{N}{2}-m^{\prime}}r_{\alpha,k_{2}}(w)^{-\frac{N}{2}}r_{\alpha,k_{3}}(w)^{\frac{N}{2}}\\ \times F_{\beta,k_{1}}(w)F_{\alpha,k_{2}}(w)F_{1,N}(w)F_{\alpha,k_{3}}(z)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ =\sum_{k=1,2}\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}z\text{d}w}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,k}(w)^{\frac{N}{2}-m^{\prime}}r_{2,N}(w)r_{\alpha,1}(w)^{\frac{N}{2}}r_{\beta,1}(w)^{\frac{N}{2}}\\ \times F_{\beta,k}(w)F_{\alpha,1}(w)F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\sum_{k=1,2}\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}z\text{d}w}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,k}(w)^{\frac{N}{2}-m^{\prime}}r_{\alpha,1}(w)^{\frac{N}{2}}r_{2,N}(w)r_{\beta,1}(w)^{\frac{N}{2}}\\ \times F_{\beta,k}(w)F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\$
@@
(53) $\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}z\text{d}w}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{N-m^{\prime}}r_{2,N}(w)r_{\alpha,1}(w)^{\frac{N}{2}}F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\$
@@
(54) $\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}z\text{d}w}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
@@
$\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,1}(w)F_{\alpha,1}(w)=F_{\beta,1}(w)+\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,1}(w)F_{\alpha,1}(w)F_{\beta,2}(w).$
To prove this identity one can multiply the left hand side on the right by
$I=F_{\beta,1}(w)+F_{\beta,2}(w).$
@@ 4.1. Preliminaries of Saddle Functions
We know that $\psi_{k}(z)$ always has at least one saddle point between the branch cuts on the negative real axis. We denote this saddle point $z_{k}^{*}$. [^8] This is the same notation used in [DK21]. We start by stating,
@@
This is largely a subset of Lemma 6.8 in [DK21]. There are a few key differences we take the time to acknowledge and justify. Firstly, since there is an overall sign change between their saddle function and the saddle function presented here, there is a reversal of maxima and minima. Additionally, while the asymptotic coordinate is restricted to the smooth region in the proof presented by Duits and Kuijlaars, we address cases where this is not necessarily true. We present the proof of these additional part below.
@@
Since we are restricting our analysis to $x<\nicefrac{{1}}{{2}}$ we can make some additional remarks regarding the general location of the saddle points. Under this restriction, $\psi_{1}(z)$ always has a single saddle point, while $\psi_{2}(z)$ has three saddle points (counting multiplicity). This is mostly a result of Lemma 4.2 along with that $\Psi^{\prime}(z)\,dz$ is a single-valued meromorphic function (for fixed $(x,y)$) on the Riemann surface $\mathcal{R}$ described before Definition 2.1. We will refer to the other two saddle points of $\psi_{2}(z)$ as $z_{3}^{*}$ and $z_{4}^{*}$. When $(x,y)\in\mathfrak{S}$, these additional saddle points also lie on the interval $(-\alpha^{-2},-\alpha^{2})$. We adopt the convention, $z_{3}^{*}<z_{2}^{*}<z_{4}^{*}$. The points $z_{3}^{*}$ and $z_{4}^{*}$ are both local minimums of $\operatorname{Re}\psi_{2}(z)$.
The contour of steepest ascent from $z_{k}^{*}$ is the contour through $z_{k}^{*}$ where $\operatorname{Im}\psi_{k}(z)$ is held constant and $\operatorname{Re}\psi_{k}(z)$ increases the fastest as we move away from the saddle point. Since $z_{1}^{*}$ is a maximum of $\operatorname{Re}\psi_{1}(z)$, the path of steepest ascent is perpendicular to the real axis at $z_{1}^{*}$ (the same is true for $z_{2}^{*}$ when the asymptotic coordinate is in the smooth region). We define the following regions of the complex plane, [^9] While this notation is the same as in [DK21] the regions will be reversed in our notation due to the change in sign of the saddle functions presented here.
@@
    The steepest ascent contour of $\psi_{1}(z;x,y)$ through $z_{1}^{*}$, denoted $\gamma_{s,1}$, is a simple closed curve that contains the interval $\alpha^{2},0]$ and intersects the positive real line at $z=1$.
@@
[…]
@@
    $\Omega^{-}_{1}$ is an open set with an unbounded component that contains a contour going around $(-\infty,-\alpha^{-2}]$ and with a bounded component that contains a contour around $\alpha^{2},0]$.
@@
    The steepest ascent contour of $\psi_{2}(z;x,y)$ through $z_{2}^{*}$, denoted $\gamma_{s,2}$, is a simple closed curve that contains the interval $\alpha^{2},0]$ and intersects the positive real line at $z=1$.
@@
[…]
@@
    $\Omega^{-}_{2}$ is an open set with an unbounded component that contains a contour going around $(-\infty,-\alpha^{-2}]$ and with a bounded component that contains a contour around $\alpha^{2},0]$.
Figure 10 shows an example of the regions $\Omega_{1}^{+}$ and $\Omega_{1}^{-}$ along with the steepest ascent contour. This lemma is a slightly modified version of Lemma 6.9 of [DK21]. The reader should refer to their text for a the proof. One can check that the proof is valid for $\psi_{1}(z)$ for any $x<\nicefrac{{1}}{{2}}$. This is because $z_{1}^{*}$ is always a local maximum of $\operatorname{Re}\psi_{1}(z)$. In the first part, we conclude that $\Omega^{+}_{1}$ is a single component, because there is exactly one saddle point on the interval $(-\alpha^{-2},-\alpha^{2})$. The case of three components is only possible if there are three saddle points on the interval $(-\alpha^{-2},-\alpha^{2})$.
@@
Let $r_{1}$ and $r_{2}$ be negative real numbers such that $r_{1}<z_{1}^{*}<r_{2}$. Then there exists contours surrounding the intervals $(-\infty,r_{1}]$ and $r_{2},0]$ that are contained entirely inside of $\Omega_{j}^{-}$ for j=1, 2j=1,\,2.
@@
The above information is enough for the analysis of $I_{2,1}^{\alpha}$ and for the analysis of $I^{\alpha}_{2,2}$ in the smooth region. In order to analyze $I^{\alpha}_{2,2}$ in the rough region, we will need some information on the contours of steepest descent and ascent in the rough region. For $x\in(0,\nicefrac{{1}}{{2}})$ and $(x,y)\in\mathfrak{R}$, the additional saddle points of $\psi_{2}(z)$, $z_{3}^{*}$ and $z_{4}^{*}$, are complex. In particular, $z_{3}^{*}=\overline{z_{4}^{*}}$. We will let $z_{3}^{*}$ denote the complex saddle point with positive imaginary part. We define the following regions of the complex plane,
@@
    $\Pi^{+}$ contains at least two and at most three bounded components. One component will always contain $z=1$, while the other guaranteed component contains at least one of the branch points, $\alpha^{2}$ or $\alpha^{-2}$. The third component, if it exists, must contain the other branch point, intersect the branch cut, and be positive distance away from the other components.